2. An ordinary die is rolled once. Find the probability that

i. An even number occurs

ii. A number greater than 4 occurs



1
Expert's answer
2021-05-26T02:31:47-0400

Solution:

"S=\{1,2,3,4,5,6\}"

(i) Let E be the event of getting an even number.

"E=\{2,4,6\}"

"P(E)=\dfrac{3}{6}=\dfrac{1}{2}"

(ii) Let G be the event of getting a number greater than 4.

"G=\{5,6\}"

"P(G)=\dfrac{2}{6}=\dfrac{1}{3}"